A grandfather clock is controlled by a swinging brass pendulum that is 1.2 m long at a temperature of 30°C.

(a) What is the length of the pendulum rod when the temperature drops to 0.0°C? (Round your answer to four significant figures.)
Answer #1

Here, we can use linear expansion equation which is given as

\DeltaL = Lo\alpha\Delta T

where Lo is original length

\alpha is coefficient of linear expansion, for Brass, it is 19e-6 / oC

Lf - Lo = Lo\alpha\Delta T

Lf = Lo\alpha\Delta T + Lo

Lf = 1.2 * 19e-6 * ( 0 - 30) + 1.2

Lf = 1.199316 m

or

Lf = 1199.316 mm